5.2 tons of pig ovaries were required to extract 16.0 mg of estrone. Assuming complete isolation of the hormones was achieved, calculate the mass percentage of estrone in the original tissue.
Answer #1

Mass of original tissue (pig ovaries) = 5.2 tons x 1000kg/ton

= 5200 kg

Mass of estrone = 16.0 mg x 1g/1000mg x 1kg/1000g

= 16*10^-6 kg

isolation of the hormones achieved = 100%

mass % of estrone in the original tissue

= Mass of estrone * 100 / Mass of original tissue

= 16*10^-6 kg * 100 / 5200 kg

= 3.08 * 10^-7 %
Answer #2

Convert the mass of pig ovaries to milligrams (mg) :

    • 1 ton = 1,000,000 grams

    • 1 gram = 1,000 mg

    • So, 5.2 tons = 5.2 × 1,000,000 × 1,000 = 5,200,000,000 mg


  1. Identify the mass of estrone extracted :

    • Estrone obtained = 16.0 mg


  2. Calculate the mass percentage
    The formula is:

    Mass percentage=(Mass of estroneMass of ovaries)×100%

    Plugging in the numbers:

    Mass percentage=(16.0mg5,200,000,000mg)×100%3.08×107%


Final Answer:
The mass percentage of estrone in the pig ovaries is approximately 0.000000308%.


source: Self
answered by: Harshwardhan kunal